YoursTruly:

Multiply and simplify: 4y(y2+3y-12)

demolisher1224:

"y2" was replaced by "y^2". 1 • -12 = -12 -12 + 1 = -11 -6 + 2 = -4 -4 + 3 = -1 -3 + 4 = 1 -2 + 6 = 4 -1 + 12 = 11 4y • (y2 + 3y - 12)

JustSaiyan:

The answer will be \[4y^3+12y^2-48y\] \[(4y)(y^2+3y+−12)\] \[(4y)(y^2)+(4y)(3y)+(4y)(−12)\]

JustSaiyan:

Even I can do it Molly.

YoursTruly:

I was in Consumer Math last semester! Don't judge me.

JustSaiyan:

All you do is simplify it. That is the only step.

YoursTruly:

I don't remember how

JustSaiyan:

Then use a calculator online, they have a million of them.

YoursTruly:

Don't tell me what to do. I'm older than you, mister. XD